# Articulate AnyMesh: Open-Vocabulary 3D Articulated Objects Modeling

### v1

@@ 4 Experiments
Implementation In movable part segmentation stage, we use Partslip++ (Zhou et al., 2023) for segmentation and replace the grounding model with DINO-X to achieve better performance. In addition, GPT4o is prompted to extract part labels of the input mesh and some of its joint configurations like joint type, parent link, etc. In articulation estimation stage, we visually prompt GPT4o for the rest of the joint configurations. For refinement stage, we implement a multi-part DMTet (Shen et al., 2021) for articulated objects representation. Each part of the object is assigned an independent DMTet network for describing its geometry and texture features. We exploit the normal-depth diffusion model and albedo-diffusion model proposed in Richdreamer (Qiu et al., 2024) for geometry and texture prior.
Baselines To the best of our knowledge, this is the first work to address the challenge of converting meshes into articulated objects in an open-vocabulary manner. Consequently, no existing baseline directly process the same input and output as Articulate AnyMesh. Instead, we evaluate individual components separately, comparing to corresponding baselines. The articulation estimation step in our pipeline takes object parts as input and outputs their joint configurations. NAP (Lei et al., 2023) and CAGE (Liu et al., 2024c), which learn diffusion models to generate articulated objects, can also generate joint configurations conditioned on object parts, making them the closest baselines for this step. Additionally, methods such as ANCSH (Li et al., 2020), OPD (Jiang et al., 2022a), and OPDmulti (Sun et al., 2024) directly estimate articulation parameters. These methods take a single observation (e.g., RGB, depth, or both) as input and simultaneously predict part segmentation and joint parameters, serving as additional baselines. We compare these methods to the combined 3D segmentation and articulation estimation steps in Articulate AnyMesh. For the refinement step, since our approach is based on Richdreamer (Qiu et al., 2024), it serves as a natural baseline for this stage.
@@
For the refinement step, since no ideal metric directly evaluates the visual quality and structural correctness of 3D objects without a reference set, we instead render 2D images of articulated objects in different states and assess these images using metrics that measure image-caption similarity. The underlying assumption is that a structurally accurate and visually plausible articulated object will produce rendered images that better correspond to its category, resulting in higher image-text correspondence scores. Specifically, we use CLIPscore (Hessel et al., 2021) and VQAscore (Lin et al., 2024) as the metrics for this evaluation.
@@ 4.1 Quantitative Experiments
Articulation estimation
As outlined in the baseline section, we evaluate the performance of articulation parameter estimation under two settings:
  • Object part mesh as input. For articulation parameter estimation using object part meshes, we compare Articulate AnyMesh against two generative diffusion models, NAP and CAGE, which represent articulated objects as graphs. In this experiment, the ground-truth shapes of all parts are provided, and the task is to estimate their articulation parameters. NAP generates joint parameters (edges) conditioned on ground-truth vertices, such as part bounding boxes, spatial locations, and shape latents. For CAGE, attributes like bounding boxes, joint types, and semantic labels are provided for each node, while joint axes and ranges are generated. We conduct this evaluation using articulated objects from PartNet-Mobility, following the train-test split used in CAGE and retraining NAP on the same split. We assess in-domain performance using objects from the test set and generalization performance on categories excluded from training. As shown in Table 1, while Articulate AnyMesh achieves comparable performance to NAP and CAGE in-domain, it significantly outperforms them in unseen categories.
@@
  • Single Observation as input. We also compare Articulate AnyMesh against three methods—ANCSH, OPD, and OPDmulti—for articulation parameter estimation from single observations. ANCSH uses a single-view point cloud as input, while OPD and OPDmulti take an RGB image (optionally with depth information) as input. These methods output part segmentation and joint parameters. To adapt Articulate AnyMesh for this setting, we modify the pipeline such that only one image (the input observation) is segmented. The resulting segmentation is projected into a point cloud with labels, which is then processed in the articulation estimation step. The results, presented in Table 2, show that Articulate AnyMesh outperforms ANCSH, OPD, and OPDmulti within their respective training domains, despite being designed as an training free, open-vocabulary method.
@@
Figure 3: Visualization of articulated objects output by the refinement step. (a) Diverse articulated objects produced by Articulate AnyMesh (the input meshes are generated with InstantMesh (Xu et al., 2024)); (b) Distinct textures created by the refinement step for the same input, with geometry and articulation parameters preserved (highlighted by the boxes); (c) A fictional object produced by Articulate AnyMesh
Figure 4: Articulate AnyMesh applies to hand-crafted meshes retrieved from Objaverse
  ANCSH OPD OPDmulti Ours
angle error $\downarrow$ 6.74 10.73 9.66 5.37
position error $\downarrow$ 0.065 0.117 0.108 0.049
Refinement
For the refinement step, we evaluate the visual quality and structural correctness of refined articulated objects using the metrics described in Section Metrics. We test three configurations on the intermediate output of the articulation estimation step: (1) No Refinement step applied, (2) Refinement step applied without random transformation (same as Richdreamer) and (3) Refinement step applied with random transformation. The surface meshes used in this experiment are generated using the text-to-3D model InstantMesh (Xu et al., 2024), conditioned on text prompts corresponding to categories in PartNet-Mobility. The results, presented in table 3, show that the highest scores are achieved with refinement incorporating random transformations, highlighting the effectiveness of our proposed refinement step.
  w/o Re w/o RT Re /w RT
CLIP Score $\uparrow$ 0.7329 0.7928 0.8205
VQA Score $\uparrow$ 0.6551 0.8164 0.9376
Table 3: Evaluation of the refinement step. Re stands for refinement and RT is the abbreviation of random transformation.
4.2 Applications
Articulated object generation By integrating an exisitng 3D generation model to generate surface meshes as the input to Articulate AnyMesh, our pipeline takes advantage of the open-vocabulary 3D generation capabilities of the 3D generation model and inherits its generative paradigm. As shown in Figure 3, our extended pipeline generates high-quality articulated objects of various categories, from everyday objects to professional tools.
Annotate 3D object datasets Other than generating from scratch, our work can also start from existing artist designed meshes. For example, we annotate part segmentation and joint structures on objects from Objaverse (Deitke et al., 2024). Such objects usually have high-quality mesh and textures and also with inner structures. Thus, we only execute the 3D segmentation part and articulation estimation part of our pipeline to annotate these objects. The results are demonstrated in Figure 4. Objects in the teaser (Figure 1) are also produced in this way.
Real-to-Sim-to-Real In this experiment, we first reconstruct the 3D surface mesh of real-world objects using 2DGS (Huang et al., 2024). We then apply Articulate AnyMesh to convert the reconstructed mesh into an articulated object represented in URDF format, making it compatible with simulation environments. Next, we sample action targets and use motion planning (Sucan et al., 2012) to avoid collisions, generating a trajectory to successfully complete the task in simulation. Finally, we replicate the trajectory in the real world and observe that the robot arm can effectively execute the task.
We focus on three tasks involving different articulated objects: pushing a drill trigger, opening a microwave door, and rotating a steering wheel, as shown in Figure 5.
Our results show that the generated trajectories can be successfully transferred to the real world, demonstrating minimal error in part segmentation and joint prediction.
Policy learning in simulation We follow DexArt (Bao et al., 2023) for articulated object manipulation policy learning and evaluate two tasks: opening a laptop lid and lifting a bucket. To augment the training set for the ”bucket” and ”laptop” experiments, we include additional articulated objects of the same categories generated by Articulate AnyMesh. Additionally, we diversify the testing set by incorporating some generated articulated objects. We fine-tune the checkpoints trained exclusively on the original dataset and evaluate their performance on the augmented testing set. The results, presented in Table 4, demonstrate that the data generated by Articulate AnyMesh effectively enhances robot learning. The articulated objects generated by Articulate AnyMeshare sourced from Objaverse. For visualizations of the data used, please refer to Appendix Section B.
  Laptop Bucket
original augment original augment
sucess rate $\uparrow$ 0.408 0.513 0.339 0.479
Table 4: Success rates of manipulation policy learning experiments, comparing policies fine-tuned on the original training set with those fine-tuned on the augmented training set. Results are reported for evaluations on the augmented test set.
5 Conclusion
Limitations and Future work Although our method creates articulated objects with semantically correct parts, the articulation parameters are not always accurate or physically grounded. This limitation arises from three key challenges: (1) existing and generated 3D meshes often lack correct physical structures, (2) current 3D segmentation methods occasionally produce incorrect results, and (3) GPT4o exhibits biases in certain scenarios. Additionally, some uncommon cases are not covered by our method—for example, when a drawer is designed to be pulled out up-front.
Appendix A Details of the Refinement Step
The refinement step builds on Richdreamer (Qiu et al., 2024) because it separately optimizes geometry and texture using two distinct diffusion models: a normal-depth diffusion model and an albedo diffusion model. Unlike computing the SDS loss purely in RGB space, which is suboptimal for geometry optimization, the normal-depth diffusion model demonstrates greater geometry awareness.
In Richdreamer, the Score Distillation Sampling process proposed by DreamFusion (Poole et al., 2022) is formulated as follows:
Given a 3D representation $\phi$, and a differentiable renderer $g$, the rendered image is $x=g(\phi)$. The SDS loss is then used to optimize the 3D representation $\phi$:
$\nabla_{\phi}\mathcal{L}_{\text{SDS}}(\phi,x=g(\phi))=\mathbb{E}_{t,\epsilon}\left[w(t)(\epsilon_{\theta}(z_{t};y,t)-\epsilon)\frac{\partial x}{\partial\phi}\right]$,
where $z_{t}$ is the noisy latent code, $\epsilon$ is the injected noise and $\epsilon_{\theta}$ is the noise predicted by a denoising model $\theta$, conditioned on timestep $t$ and text embedding $y$. The term $w(t)$ is a timestep dependent weighting factor.
In Richdreamer and other previous works the 3D representation $\phi$ is static, whereas in our case, it is articulated. Omitting other attributes, we denote the 3D representation of articulated objects as $\phi_{q}$, where $q$ is a vector representing joint positions. During optimization, the base of the articulated object remains fixed. Since non-fixed joints of interest (revolute, prismatic, and continuous) all have one degree of freedom, each element in $q$ corresponds to the position of a non-fixed joint in $\phi_{q}$. The articulated object in its rest configuration is denoted as $\phi_{q_{0}}$. A transformation function $T$ maps $\phi_{q_{0}}$ to $\phi_{q}=T(\phi_{q_{0}},q)$ given the desired joint positions $q$.
The optimization process for Richdreamer (Qiu et al., 2024) and other previous SDS optimization methods can be briefly summarized by the following pseudo-code:
FOR i IN iterations:
    render an image x (x = g(phi))
    sample timestep t
    compute SDS loss
    update optimizer
In our approach, we extend this process by sampling joint positions and transforming object parts accordingly:
FOR i IN iterations:
    sample joint position q
    transform parts according to q (phi_q = T(phi_q0, q))
    render an image x (x = g(phi_q))
    sample timestep t
    compute SDS loss
    update optimizer
We also provide visualizations of the input and output of the refinement step, as shown in Figure 6. The pencil case and lighter in the second and third rows have manually set joint parameters and shape primitives as link geometries. The refinement step successfully optimizes storage space for the drawer and pencil case and generates a nozzle structure for the lighter, and produces plausible textures.
Figure 6: Visualization of the input and output of the refinement step.
Appendix B Implementation of policy learning experiment
@@
In this experiment, we evaluate two tasks proposed by DexArt (Bao et al., 2023). The first task is lifting a bucket, where a dexterous hand must grasp the bucket handle and lift it. In the original implementation, the bucket handle is thickened to prevent severe penetration, and we adhere to this setting. The second task is opening a laptop lid, where a dexterous hand must open the lid to a 90-degree angle.
@@
For both tasks, we fine-tune the ”no-pretrain” policy checkpoint provided by DexArt’s official GitHub repository using a combined training set consisting of DexArt’s original samples (sourced from PartNet-Mobility (Xiang et al., 2020)) and additional samples (sourced from Objaverse (Deitke et al., 2024)) generated by Articulate AnyMesh. As a baseline, we fine-tune the policy for the same number of steps but only use the original dataset. During evaluation, the new policies are tested on both the original test set and the additional test samples generated by Articulate AnyMesh. Visualizations of the datasets used are provided in Figures 7, 8, and 9.
@@
Figure 7: Visualization of the original training set for the bucket task from DexArt (Bao et al., 2023) and the augmented training set collected using Articulate AnyMesh.
@@
Figure 8: Visualization of the original training set for the laptop task from DexArt (Bao et al., 2023) and the augmented training set collected using Articulate AnyMesh.
@@
Figure 9: Visualization of the original test set used by DexArt (Bao et al., 2023) and augmented test set collected using Articulate AnyMesh.
Appendix C Prompting details
The prompt for recognizing parts in an rendered image of a input mesh:
You have a good understanding of the structure of articulated objects. Your job is to assist the user to analyze the structure of an object. Specifically, the user will give you an image of an articulated object, and your task is to recognize the main parts of that object. You should give your answer in the following format:
‘‘‘part_list
(1) part_name: name of the part; description: a brief description about the part, and how it moves
(2) part_name: name of the part; description: a brief description about the part, and how it moves
Remember:
(1) Do not answer anything not asked.
(2) Your answer should be purely based on the input image, do not imagine anything.
(3) If there are multiple parts with the same semantic, just add one part to the list. For example, if there are four wheels, just add one part whose name is wheel.
The prompt to generate the link and joint configuration of the articulated object, and exclude non-movable parts:
SYSTEM PROMPT:
You have a good understanding of the structure of articulated objects. You are very familiar with URDF format. Your job is to assist the user to analyze the structure of an articulated object. Specifically, the user will name an object and then give you the main parts of that object. You will have to group these parts into links and then give the joints connecting these links. You should give your answer in the following format:
‘‘‘articulation tree
parts:
(1) part_name: name of the recognized part;
links:
(1) link_name: name of the link;
joints:
(1) joint_name: name of the joint; joint_type: type of the joint; parent_link: name of the parent link; child_link: name of the child link; joint_limit: [lower limit, upper limit];
For example:
‘‘‘articulation tree
parts:
(1) part_name: Front windshield;
(2) part_name: Doors;
(3) part_name: Headlights;
(4) part_name: Wheels;
(5) part_name: Windows;
links:
(1) link_name: Chasis;
(2) link_name: Doors;
(3) link_name: Wheels;
(4) link_name: Windows;
joints:
(1) joint_name: wheel_chasis_joint; joint_type: continuous; parent_link: Chasis ; child_link: Wheels; joint_limit: None;
(2) joint_name: door_chasis_joint; joint_type: revolute; parent_link: Chasis ; child_link: Doors; joint_limit: [0, 90];
(3) joint_name: door_chasis_joint; joint_type: prismatic; parent_link: Chasis ; child_link: Windows; joint_limit: [0, 1];
Remember:
(1) Do not answer anything not asked.
(2) If a part is actually movable in any direction, its joint type is floating.
(3) Available joint types are: fixed, prismatic, revolute, continuous and floating.
(4) For joint_type, only answer one word (among the available types), do not answer anything else.
(5) For every part that is not fixed, there must be a unique link for it.
(6) For parts that are fixed, try to group as many as you can.
(7) Joint limit must be given for prismatic joints and revolute joints. The unit of a revolute joint limit is degrees. The unit of a prismatic joint limit is the size of its child link in its translation direction. The joint limit should be two numbers.
USER PROMPT:
Object: OBJECT_NAME
Parts: RECOGNIZED_PARTS
The following prompts are for joint parameter estimation:
For revolute joints, we first prompt GPT4o to determine whether both points of the joint are on the surface:
You are an assistant with a deep understanding of the structure of objects. Your task is to help users determine the hinge position of some parts of a given object mesh using common sense. It is important to note that the object is represented by a mesh, so you only have access to the objects surface and no access to its inner structure. The term "hinge" here does not only refer to the mechanical structure of a hinge, but also has a broader meaning. For example, the connection between a cardboard box lid and the body of the box is also considered a hinge.
Specifically, the user will provide you with an object, and the part for which the hinge position needs to be predicted will be specified. You will have to decide whether (1) both ends of the hinge are positioned on the surface of the object or (2) only one end of the hinge is positioned near the surface of the object, and the other end is inside the object.
For example, the hinge of a door has both its ends positioned on the door frame, which is recognizable and falls into the first category. The hinge of a wheel has one end recognizable on the center of the wheel, and its other end hidden inside the car (normally an object mesh of a car will not have detailed mechanical structures), which falls into the second category.
Please give your answer in the following format:
‘‘‘hinge_info
description: a brief description of the hinge
choice: (1) or (2)
Based on GPT4o’s selection, we provide a tailored prompt. For choice (1), the specific prompt is as follows. Once GPT4o selects two or more points, we fit a line through these points, defining the rotation axis.
You are an assistant with a deep understanding of the structure of objects. Your task is to answer some questions about the input image of an object. The input image is of a {object_name}. The image has some points marked, each with an numerical ID as a label. Please select the points that are on the rotation axis of the {part_name} of the {object_name}. Give your answer in the following format:
‘‘‘hinge points
description: a brief description of the location of the rotation axis and the selected points
selected IDs: ID of selected point1, ID of selected point2, ... (for example 1,3)
Remember:
(1) Do not answer anything not asked for.
(2) Select two or more points.
(3) Give your answer based on the provided image.
For choice (2), the prompt is as follows. The rotation axis is determined using the point selected by GPT4o and the normal of the connected area.
You are an assistant with a deep understanding of the structure of objects. Your task is to answer some questions about the input image of an object. The input image is of a {object_name}. The image has some points marked, each with an numerical ID as a label. Please select the point that is on the rotation axis of the {part_name} of the {object_name}. Give your answer in the following format:
‘‘‘hinge points
description: a brief description of the location of the rotation axis and the selected points
selected IDs: ID of the selected point
Remember:
(1) Do not answer anything not asked for.
(2) Only select one point that is the most suitable.
(3) Give your answer based on the provided image.
For prismatic joints, we prompt GPT4o to determine whether its child link moves in and out or along the surface:
You are an assistant with a deep understanding of the structure of objects. Your task is to help users to determine the translation direction of some parts of a given object mesh using common sense. It is important to note that the object is represented by a mesh, so you only have access to the objects surface and no access to its inner structure.
Specifically, the user will provide you with an object and the part for which the translation direction needs to be predicted will be specified. You will have to decide whether the translation direction is outwards from/inwards towards the mesh, or along the surface of the mesh.
When a part moves outwards, you will see more of that part coming out from the object. When a part moves inwards, you will see portions of that part going into the object. When a part moves along the surface of an object, you still see the exact same part.
For example, a pressing button can be pressed inwards (when you press it, the button goes into the object), the telescopic handle of a suitcase can be pulled outwards (when you pull it, the entire handle comes out of the suitcase), and a stick shift moves along the surface of a shift pattern (when you are shifting, the shift does not go into the transmission or out of the transmission).
Please give your answer in the following format:
‘‘‘translation_axis_info
description: a brief description of the translation axis
choice: outward/inward or surface
If the child link moves inward or outward, the translation direction is defined by the normal vector of the connected area. If it moves along the surface, we draw four arrows (originating from the child link’s center and pointing up, down, left, and right) on the image and prompt GPT4o to select the translation direction:
You are an assistant with a deep understanding of the structure of objects. Your task is to answer some questions about the input image of an object. The input image is of a {object_name}. The image has some arrows marked, each with a different color. Please select the arrow that indicate the translation direction of the {part_name} of the {object_name}. Give your answer in the following format:
‘‘‘sliding direction
description: a brief description of the direction of the sliding axis and the selected arrow
selected arrow: color of the arrow (in red, yellow, blue, green)
Remember:
(1) Do not answer anything not asked for.
(2) Select one arrow.
The selected arrow is then projected onto the plane fitted to the connecting area, and the translation direction is defined by the direction of the projected arrow.
4. Validate joint limits. For revolute joints, we incrementally rotate the child link based on the joint parameters and identify the maximum range where penetration remains below a predefined threshold. For prismatic joints, we follow the joint limits provided by GPT4o.

### v2

@@ 4 Experiments
Implementation In the movable part segmentation stage, we use Partslip++[40] for segmentation and replace the grounding model with a pipeline that uses set-of-mark visual prompting[64] to enable the VLM to perform grounding for better performance. In addition, GPT4o is prompted to extract part labels of the input mesh and some of its joint configurations like joint type, parent link, joint limit, etc. In articulation estimation stage, we visually prompt GPT4o for the rest of the joint configurations. In the Post-Processing stage, we employ HoloPart [46] for 3D shape completion and Meshy [47] for texture generation.
Baselines To the best of our knowledge, this is the first work to tackle the challenge of converting meshes into articulated objects in an open-vocabulary manner. As a result, no existing baseline directly processes the same input and output as Articulate AnyMesh. Therefore, we compare against prior works that perform joint parameter estimation only on object categories seen during training.
URDFormer [2] takes a front-view image as input and sequentially outputs part segmentation, joint parameters, and part geometry. Real2Code [29] takes multi-view images as input and performs part segmentation, shape completion, and joint parameter estimation through code generation. For these two methods, we compare Articulate AnyMeshon the PartNet-Mobility dataset [33]. Additional comparisons with estimation methods such as ANCSH[44], OPD[52], and OPD-Multi[53], as well as generative methods like NAP[27] and CAGE[28], are provided in Appendix section A. We also provide an ablation study analyzing the impact of different vision-language models (VLMs), including GPT, Gemini, and Claude, in Appendix section B.
@@ 4.1 Quantitative Experiments
We compare Articulate AnyMeshto URDFormer and Real2Code on the PartNet-Mobility dataset. To ensure a comprehensive evaluation, we assess joint estimation performance on both in-domain (ID) and out-of-domain (OOD) object categories. The in-domain categories include the five object types used to train URDFormer: oven, fridge, cabinet, washer, and dishwasher. For Real2Code, we fine-tune a CodeLlama model on objects from the same categories. We align the inputs of all methods as closely as possible to ensure a fair comparison: URDFormer takes a front-view image and is additionally provided with ground-truth part bounding boxes to predict joint parameters per part; Real2Code takes ground-truth part meshes and performs joint estimation via code generation; Articulate AnyMeshtakes the ground-truth mesh, performs part segmentation, and estimates joint parameters via visual prompting. Results in table 2 show that although URDFormer and Real2Code perform just slightly worse on in-domain categories, their performance drops significantly on out-of-domain objects. In contrast, Articulate AnyMeshmaintains low error across both ID and OOD settings. This highlights the key advantage of our approach: it is not constrained by the biases of existing articulated object datasets and generalizes well beyond them.
Figure 3: Results of articulated object generation. Each example shows (from left to right): the raw surface mesh generated by Meshy, the completed geometry, and the textured object, all rendered under two different poses.
Figure 4: Results of Articulate AnyMesh applied to hand-crafted meshes retrieved from Objaverse.
  ID OOD
$\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$
URDFormer 13.674 0.144 33.864 0.299
Real2Code 7.950 0.167 24.035 0.194
Ours 6.257 0.076 6.582 0.083
Table 1: Comparison of joint estimation accuracy—including angle and position error—between Articulate AnyMesh, URDFormer and Real2Code.
  Laptop Bucket
original aug original aug
succ. rate $\uparrow$ 0.408 0.513 0.339 0.479
Table 2: Success rates of manipulation, comparing fine-tuning on original vs. augmented training sets. Evaluated on the augmented test set.
4.2 Applications
Articulated object generation By integrating an existing 3D generation model (Meshy[47]) to generate surface meshes as inputs to Articulate AnyMesh, our pipeline exploits the open-vocabulary 3D generation capabilities of the 3D generation model and inherits its generative paradigm. As in Figure 3, our extended pipeline generates high-quality articulated objects of various categories.
Annotate 3D object datasets Other than generating from scratch, Articulate AnyMeshcan also start from existing artist designed meshes. For example, we annotate part segmentation and joint structures on objects from Objaverse [1]. Such objects usually have high-quality mesh and textures and also with inner structures. Thus, we only execute the 3D segmentation part and articulation estimation part of our pipeline to annotate these objects. The results are demonstrated in Figure 4. Objects in the teaser (Figure 1) are also produced in this way.
Real-to-Sim-to-Real In this experiment, we first reconstruct the 3D surface mesh of real-world objects using 2DGS [65]. We then apply Articulate AnyMesh to convert the reconstructed mesh into an articulated object represented in URDF format, making it compatible with simulation environments. Next, we sample action targets and use motion planning [66] to avoid collisions, generating a trajectory to successfully complete the task in simulation. Finally, we replicate the trajectory in the real world and observe that the robot arm can effectively execute the task.
We focus on three tasks involving different articulated objects: pushing a drill trigger, opening a microwave door, and rotating a steering wheel, as shown in Figure 5. Our results show that the generated trajectories can be successfully transferred to the real world, demonstrating minimal error in part segmentation and joint prediction.
Policy learning in simulation We follow DexArt [67] for articulated object manipulation policy learning and evaluate two tasks: opening a laptop lid and lifting a bucket. To augment the training set for the ”bucket” and ”laptop” experiments, we include additional articulated objects of the same categories generated by Articulate AnyMesh. Additionally, we diversify the testing set by incorporating some generated articulated objects. We fine-tune the checkpoints trained exclusively on the original dataset and evaluate their performance on the augmented testing set. The results, presented in Table 2, demonstrate that the data generated by Articulate AnyMesh effectively enhances robot learning. The articulated objects generated by Articulate AnyMeshare sourced from Objaverse. For visualizations of the data used and implementation details, please refer to Appendix Section C.
5 Conclusion
6 Limitations
Although our method creates articulated objects with semantically correct parts, the articulation parameters are not always accurate or physically grounded. This limitation arises from three key challenges: (1) existing and generated 3D meshes often lack correct physical structures, (2) current 3D segmentation methods occasionally produce incorrect results, and (3) GPT4o exhibits biases in certain scenarios. Additionally, some uncommon cases are not covered by our method—for example, when a drawer is designed to be pulled out up-front.
Appendix A More quantitative results on joint parameter estimation
In this section, we provide additional comparisons with methods not covered in the main paper for joint parameter estimation. Specifically, we include two generative approaches—NAP and CAGE—that leverage diffusion models to synthesize articulated objects and can generate joint configurations conditioned on object parts. In addition to these generative baselines, we also compare against discriminative methods such as ANCSH[44], OPD[52], and OPDmulti[53]. These methods take a single observation (e.g., RGB, depth, or both) as input and jointly predict part segmentation and articulation parameters. Together, these methods provide a broader context for evaluating joint estimation performance.
@@ Appendix A More quantitative results on joint parameter estimation
We first evaluate performance in the single-observation setting, using the OPD-Synth dataset. In this experiment, we compare Articulate AnyMeshagainst ANCSH, OPD, and OPDmulti for articulation parameter estimation from a single-view input. ANCSH uses a single-view point cloud as input, while OPD and OPDmulti operate on RGB images, optionally supplemented with depth. All three methods jointly predict part segmentation and joint parameters. To ensure a fair comparison, we adapt Articulate AnyMeshto this setting by restricting segmentation to a single image, which is then projected into a point cloud with part labels for articulation parameter estimation. As reported in Table 3, Articulate AnyMesh consistently outperforms all three methods within their respective training domains, despite being designed as a training-free, open-vocabulary approach.
We also evaluate articulation parameter estimation using ground-truth shapes of object parts as input, from PartNet-Mobility. Here, we compare Articulate AnyMeshwith the generative diffusion-based models NAP and CAGE, both of which represent articulated objects as part graphs. NAP predicts joint parameters conditioned on part-level geometry descriptors such as bounding boxes, spatial locations, and learned shape latents. CAGE, by contrast, generates joint axes and ranges given attributes including part bounding boxes, semantic labels, and joint types. Following the experimental setup in CAGE, we use the PartNet-Mobility dataset and retrain NAP on the same train-test split. We evaluate both in-domain performance on test categories and generalization to held-out categories. As shown in Table4, while all three methods perform comparably in-domain, Articulate AnyMesh substantially outperforms NAP and CAGE on previously unseen object categories, demonstrating superior generalization.
  ANCSH OPD OPDmulti Ours
angle error $\downarrow$ 6.74 10.73 9.66 5.37
position error $\downarrow$ 0.065 0.117 0.108 0.049
@@
Appendix B Ablation study of VLM
  ID OOD
$\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$
URDFormer 13.674 0.144 33.864 0.299
Real2Code 7.950 0.167 24.035 0.194
Ours (Claude) 9.291 0.160 7.541 0.096
Ours (Gemini) 6.477 0.239 6.612 0.081
Ours (GPT) 6.257 0.076 6.582 0.083
Table 5: Comparison of joint estimation accuracy (angle and position error) using different vision-language models.
To investigate the effect of different vision-language model (VLM) choices, we conducted joint parameter estimation experiments using three VLMs: GPT-4o, Claude 3.7 Sonnet, and Gemini 2.0 Flash. Ground-truth part meshes from selected PartNet-Mobility categories (consistent with the setting used in the comparison with URDFormer and Real2Code, Table 2) were used as input for part segmentation and joint prediction. The results, presented in Table 5, indicate that GPT-4o slightly outperforms Claude and Gemini, although the performance gap is relatively small. Notably, all three variants significantly outperform URDFormer and Real2Code across both in-domain and out-of-domain settings.
Appendix C Implementation of policy learning experiment
@@
In this experiment, we evaluate two tasks proposed by DexArt [67]. The first task is lifting a bucket, where a dexterous hand must grasp the bucket handle and lift it. In the original implementation, the bucket handle is thickened to prevent severe penetration, and we adhere to this setting. The second task is opening a laptop lid, where a dexterous hand must open the lid to a 90-degree angle.
@@
For both tasks, we fine-tune the ”no-pretrain” policy checkpoint provided by DexArt’s official GitHub repository using a combined training set consisting of DexArt’s original samples (sourced from PartNet-Mobility [33]) and additional samples (sourced from Objaverse [1]) generated by Articulate AnyMesh. As a baseline, we fine-tune the policy for the same number of steps but only use the original dataset. During evaluation, the new policies are tested on both the original test set and the additional test samples generated by Articulate AnyMesh. We conducted 100 evaluations on the test set, where each evaluation included all test set objects and included randomization in position and rotation. The final success rate was obtained by averaging across all evaluations. Visualizations of the datasets used are provided in Figures 6, 7, and 8.
@@
Figure 6: Visualization of the original training set for the bucket task from DexArt [67] and the augmented training set collected using Articulate AnyMesh.
@@
Figure 7: Visualization of the original training set for the laptop task from DexArt [67] and the augmented training set collected using Articulate AnyMesh.
@@
Figure 8: Visualization of the original test set used by DexArt [67] and augmented test set collected using Articulate AnyMesh.